OpenStudy (anonymous):

The legs of a right triangle are 5m and 12m long. What is the length of the hypotenuse? Explain how you found your answer.

OpenStudy (ivettef365):

use Pythagorean theorem

OpenStudy (ivettef365):

a^2 + b^2 = c^2

OpenStudy (whpalmer4):

The solution to this is one of the so-called "Pythagorean triples" — sets of \(a,b,c\) where all 3 are integers and satisfy \(a^2+b^2=c^2\). The most famous one is of course 3,4,5. This one is maybe the second most famous :-) Carpenters use the 3,4,5 triple all the time with a ruler to make square corners.

OpenStudy (anonymous):

is it 169?

OpenStudy (anonymous):

@ivettef365

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

what is the square root of 169?

OpenStudy (whpalmer4):

\[c^2 = 5^2+12^2 = 25+144 = 169\]but you need \[c = \sqrt{169}=\]

OpenStudy (anonymous):

13

OpenStudy (anonymous):

@whpalmer4

OpenStudy (ivettef365):

correct